ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 186, 2011

Albert, M. and A. Maier Additive polynomials for finite groups of Lie type

125

André, N. and I. Shafrir On a minimization problem with a mass constraint involving a potential vanishing on two curves

97

Antei, M. On the abelian fundamental group scheme of a family of varieties

427

Bekh-Ochir, C. and S. A. Rankin Examples of associative algebras for which the T -space of central polynomials is not finitely based

333

Ben-Bassat, O. and N. Solomon Cohomology groups of deformations of line bundles on complex tori

349

Crespo, T. and Z. Hajto Picard–Vessiot theory and the Jacobian problem

401

Damian, E. The generation of the augmentation ideal in profinite groups

447

Di Vincenzo, O. M. and E. Spinelli A characterization of ∗-minimal algebras with involution

381

Díaz, R., I. Garijo and R. A. Hidalgo Uniformization of conformal involutions on stable Riemann surfaces

297

Dijkstra, J. J. Characterizing stable complete Erdős space  477

Dubois, L. An explicit Berry–Esséen bound for uniformly expanding maps on the interval

221

Fokkink, R., C. Kraaikamp and H. Nakada On Schweiger’s problems on fully subtractive algorithms

285

Garijo, I. See R. Díaz Gluskin, E. and A. Olevskii Invertibility of sub-matrices and the

octahedron width theorem

61

Gonçalves, D. J. and P. Koshlukov A-identities for upper triangular matrices: a question of Henke and Regev

407

Hajto, Z. See T. Crespo Hidalgo, R. A. See R. Díaz Izhakian, Z. and L. Rowen Supertropical matrix algebra II: solving

tropical equations

69

Kazarin, L. S., A. Martínez-Pastor and M. D. Pérez-Ramos On the Sylow graph of a group and Sylow normalizers

251

Koshlukov, P. See D. J. Gonçalves Kraaikamp, C. See R. Fokkink Maier, A. See M. Albert Martínez-Pastor, A. See L. S. Kazarin Nakada, H. See R. Fokkink

Nishry, A. The hole probability for Gaussian entire functions  197

Olevskii, A. See E. Gluskin Pérez-Ramos, M. D. See L. S. Kazarin Rankin, S. A. See C. Bekh-Ochir Rowen, L. See Z. Izhakian Shafrir, I. See N. André Sivatski, A. S. Representation of polynomials as products of two val-

ues of a quadratic form

273

Solomon, N. See O. Ben-Bassat Spinelli, E. See O. M. Di Vincenzo Tanaka, N. A viability theorem of stochastic semilinear evolution

equations

1

Yomdin, Y. Remez-type inequality for discrete sets  45